OpenStudy (anonymous):

(1-(3/x))^x x-> infinity solve using l'hoptials rule

OpenStudy (loser66):

@amistre64 I don't know how to solve it. Please, help

OpenStudy (amistre64):

ln(1-3/x) --------- = ln y 1/x as x to infinity we get 0/0 right?

OpenStudy (amistre64):

ln(1-3x^-1) ----------- x^-1 3x^-2 --------------- -x^-2 (1-3x^-1) -3 -------- = -3 = ln(y) 1- 3/x but id have to check the wolf to be sure

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=limit%28x+to+inf%29+%281-3%2Fx%29%5Ex yeah, e^-3 -3 = ln(y) e^-3 = y

OpenStudy (loser66):

nope @amistre64 the problem is \[lim_{x\rightarrow \infty} (1-\dfrac{3}{x})^x\]

OpenStudy (amistre64):

thats what i did ....

OpenStudy (amistre64):

a^x = y x lna = lny lna ---- = lny 1/x etc

OpenStudy (anonymous):

im confused how you got from the first step to the second one, can you explain it to me thank you

OpenStudy (amistre64):

lhop works on indeterminate forms: 0/0

OpenStudy (loser66):

nvm , I got you

OpenStudy (amistre64):

\[3x=\frac{3}{1/x}\]

OpenStudy (amistre64):

i didnt label the steps, can you be more specfic?

OpenStudy (amistre64):

openstudy is not playing well on my end .... gotta love a broken site

OpenStudy (anonymous):

never mind i understand now thank you very much, i was confused because i didnt know you applied l'hoptials rule in your steps you were taking the derivative of hte bottom and top i was a bit confused. thank you very much!

OpenStudy (amistre64):

\[\lim_{x \to a}f^x(x)\] let f^x(x) = y, such that: \[f^x(x)=y\] \[x~ln(f(x))=ln(y)\] \[\frac{ln(f(x))}{x^{-1}}=ln(y)\]now lHop if it takes the indeterminant form \[\frac{f'(x)}{-x^{-2}~f(x)}\] when you find the limit, then reverse the ln stuff

OpenStudy (amistre64):

:) youre welcome